OpenStudy (anonymous):

What is the value of x in the solution to the following system of equations? x - y = -3 x + 3y = 5

OpenStudy (ivettef365):

isolate y on the first equation and you have y = x + 3 now substitute (x+3) on the second equation and solve for x

OpenStudy (dumbcow):

x=-1 ....others will explain :)

OpenStudy (anonymous):

-3+4y=5 4y=8 y=2

OpenStudy (anonymous):

then x-2=-3 x=-1

OpenStudy (anonymous):

dumbcow is right

OpenStudy (anonymous):

thanks